OpenStudy (anonymous):

Evaluate the integral. (Round your answer to four decimal places.) the integral of 5 csc^3(x) dx from (pi/3 to pi/6)

OpenStudy (anonymous):

\[5\int\limits_{\pi/6}^{\pi/3} \csc^3(x)\]